After a detailed analysis of sales by classification, the hosiery buyer determined a decline in the panty hose category. Last year’s sales for this category were $75,000. This year, the sales for this classification are planned at $60,000. What is the percentage of sales decline?
Answer #1

Declining in sale = 75000-60000 = $15,000

Percentage of sales decline = 15000/75000 = 20%